# ANID: How Far Are We? Evaluating the Discrepancies Between AI-synthesized Images and Natural Images through Multimodal Guidance

### v1

@@ 2. Semantic Alignment:
Generative models require the referenced prompts as semantic guidance to generate images. Therefore, semantic alignment can be an important quality indicator of the AIGIs. We use the widely used CLIP model (Radford et al. 2021b) to measure semantic alignment and compare the different CLIP Scores from natural text, image, and the paired AI-generated image-text prompt. The studied AI-generated images were originally generated from the 8 Representative generative models based on three types of prompts as described in Appendix Sec. B; the semantic alignment evaluation can provide a fine-grained analysis of the potential differences in semantic alignment with respect to both unimodal and multimodal guidance.
3. Aesthetic Appeal:
@@
Aesthetic Appeal evaluation is to estimate the visual appeal and artistic quality of images, reflecting the visual attractiveness and artistic quality of images. We utilize the classical aesthetic measures NIMA (Esfandarani and Milanfar 2018) and LAION-AES (Schuhmann et al. 2022) as the quantitative metrics and compare the aesthetic scores for the paired AI-generated and natural images.
4. Downstream Applicability:
@@
This aspect is devised to investigate the practical utility of AI-generated images in downstream tasks and evaluate whether AI-generated images can have different practical utilities with realistic natural images in downstream application tasks. We mainly focus on two classical downstream tasks, e.g., image recognition and object segmentation. For the image recognition task, we evaluate the Classification Mismatch Rate (MR) rates between AI-generated images and natural images when using a pre-trained image recognizer like ResNet-152 (He et al. 2016) model. For the object segmentation task, we use the Intersection over Union (IoU) metric (Everingham et al. 2010) as quantitative measures and investigate the segmentation results for AI-generated images with the segmentation model U2NET (Qin et al. 2020).
5. Human Assessment:
@@
We involve human assessments to coordinate the above evaluation aspects. We develop the human assessment interface to demonstrate AI-generated images with the referenced natural images and text descriptions and collect human ratings (on a scale from 1 to 5) alongside the following three aspects: naive image quality, semantic alignment, and Aesthetic Appeal as human assessments. Human participants are provided with example images before the evaluation to understand high and low scores without being given specific numerical values, ensuring unbiased and informed ratings. Detailed information about human assessment procedures is provided in Appendix Sec. D.
@@
Through the integration of diverse and comprehensive evaluation dimensions, our fine-grained evaluation framework offers a systematic assessment solution to investigate and interpret the differences that still remain between AI-generated images and natural images.
@@
Figure 2: We evaluate eight image-generating models from 6 aspects. The numerical values in the radar chart represent the mean difference rate of each model.
Evaluation
@@
RQ1: What are the fine-grained discrepancies between AIGIs and natural images?
@@
RQ2: How significant are the discrepancies in semantic alignment between AIGIs and natural images in different types of guided prompts?
@@
Experimental Setting
We use the described quantitative evaluation metrics, and details of all metrics are presented in Appendix Sec. B.
@@
For full-reference metrics like pixel-level image quality, we directly report the calculated value cause it already can present the discrepancy.
@@
Difference $\text{{Difference Rate}}=\frac{1}{n}\sum_{i=1}^{n}\frac{\left|m(\mathbf{X}_{i})-m(\mathbf{N}_{i})\right|}{m(\mathbf{N}_{i})}$ (1)
where $n$ is the number of images, $m$ represents the metric, $\mathbf{X}$ denotes the generated images, and $\mathbf{N}$ refers to the referenced natural images.
RQ1: Naive Quality Results
To evaluate the image quality of AI-generated images versus natural images, we conducted a comprehensive analysis using a variety of metrics that assess visual feature quality, naturalness, and similarity of images. Our findings reveal substantial discrepancies between AIGIs and natural images across these dimensions.
RQ1.a Pixel-level
We evaluated the visual similarity between AI-generated and natural images using structural metrics like SSIM, LPIPS, DISTS, and PSNR. The results, as shown in Table 2, indicate that AI-generated images exhibit significantly lower similarity to natural counterparts. AI-generated images show a 20% to 50% reduction in SSIM, highlighting a major loss in structural fidelity. LPIPS and DISTS further reveal notable perceptual dissimilarities, with LPIPS scores ranging from 0.19 to 0.60. Additionally, PSNR values, between 12.42 and 23.25, suggest that AI-generated images have higher noise levels, reducing their overall visual similarity.
SD_V14 SD_V15 Versatile SD_V21 SD_XL DALL·E 2
SSIM ↑ 0.35 0.35 0.46 0.44 0.70 0.38
PSNR ↑ 13.81 13.80 16.52 16.00 23.25 12.42
VIF ↑ 0.02 0.02 0.05 0.04 0.17 0.03
@@
MAD ↓ 211.61 211.73 196.86 202.29 143.66 214.86
Table 2: Full-reference Evaluation Results (I2I).
RQ1.b Frame-level
@@
We assessed the frame-level quality of AI-generated and natural images using a suite of metrics, including MUSIQ, DBCNN, HyperIQA, LIQE, BRISQUE, NRQM and NIQE. Figure 2 uses radar figures to visualize the results from the six metrics, and indicates that there are significant quality differences between AI-generated and natural images. These deviations typically range from 10% to 20%, irrespective of the guidance method used, with some instances showing quality differences as high as 30%. This underscores the considerable gaps between AIGIs and natural images in terms of global image quality. Moreover, the results of NRQM and NIQE reveal that AI-generated images significantly differ from natural images in terms of naturalness. Specifically, AI-generated images show notable deviations from natural images in visual realism and adherence to natural scene statistics. The NRQM difference rates range from 5% to 20%, while the NIQE difference rates range from 10% to 50%, highlighting the challenges AI-generated images face in replicating the inherent naturalness of real-world scenes. More detailed results with more metrics are presented in Appendix Sec. C.
@@
Figure 3: The FID and Inception Score of AI-generated images from different source models and guidance.
RQ1.c Visual Content Distribution
@@
We present the FID and Inception Score results, presented in Figure 3, suggesting that AI-generated images often exhibit relatively higher FID scores and lower Inception Scores. This finding indicates that current generative models, particularly GAN-based models and earlier versions of Stable Diffusion, struggle to consistently produce high-quality images. Notably, these models tend to generate better images when a reference image is provided (TI2I), as it has more detailed guidance.
@@
In conclusion, our evaluation of image visual feature quality, naturalness, and similarity reveals significant disparities between AI-generated and natural images. These differences highlight the current limitations of AIGC technologies in matching the visual quality and realism of natural images.
RQ2: Alignment Results
@@
In our evaluation of alignment results, we assessed how well AI-generated images correspond to their respective reference prompts across different guidance types: Text-to-Image (T2I), Image-to-Image (I2I), and Text-and-Image-to-Image (TI2I). And we find that AI-generated images often struggle to maintain high alignment with their prompts, especially the image-only reference provided. The results, shown in Table 3, indicate that AI-generated images typically achieve lower CLIP Score compared to natural images, with alignment discrepancies ranging from 4.02% to 36.69%. We also find that even if the AIGIs can get a higher CLIP score, especially with text and image-guided, they are looking stranger to humans as Table 3 illustrated. The CLIP score, which measures the alignment between text descriptions and generated images by embedding both into a shared space and calculating cosine similarity, revealed a noticeable gap in how accurately AIGC images reflect the input prompts. This suggests that current AIGC models often struggle to maintain high fidelity to the provided references, leading to variations that may impact the intended alignment. Our findings underscore the challenges that these models face in consistently generating content to accurately reflect the specified input conditions, highlighting both their strengths and limitations in real-world applications.
GALIP DFGAN SD_V14 SD_V15 Versatile SD_V21 SD_XL DALL·E 2
T2I 10.72 18.51 9.78 9.69 9.71 9.93 10.20 9.46
I2I - - 36.46 36.69 4.92 26.96 3.98 5.41
TI2I - - 9.21 9.20 8.03 8.48 4.02 -
Table 3: The difference rate of CLIP score (%).
RQ3: Aesthetic Appeal
@@
We assessed the aesthetic appeal of AI-generated and natural images using NIMA and LAION-AES metrics and found that AI-generated images generally fall short of natural images in terms of aesthetic quality. Figure 4 presents the results, which reveal that AI-generated images receive lower scores for aesthetic appeal, with discrepancies ranging from 3.34% to 23.85%. These findings indicate that while AI models can produce visually appealing images, they often lack the nuanced artistic quality and emotional impact that are characteristic of natural images.
Figure 4: The aesthetic difference rate between AI-generated and naturally generated images.
RQ4: Applicability Results
@@
When evaluating the applicability of AI-generated images, we focused on two crucial downstream tasks: image recognition and semantic segmentation. Our findings indicate significant discrepancies between the performance of AIGC images and natural images in these tasks.
@@
For image recognition, we measured the classification mismatch rate, which quantifies the differences in predicted labels between AIGC images and their natural counterparts. The results in Table 4 showed a substantial gap, with the mismatch rate ranging from 29.89% to 94.44%. This indicates that AI-generated images often fail to achieve the same level of accuracy as natural images, highlighting the limitations of current AIGC models in producing reliable content for image recognition tasks. The high mismatch rate underscores the need for further advancements in generative models to improve their applicability in real-world recognition tasks.
Overall, our results indicate that AI-generated images currently exhibit significant limitations in applicability for both image recognition and semantic segmentation tasks. The high classification mismatch rates and variable mean IoU values highlight the gap between AIGC and natural images in practical applications. These findings emphasize the importance of ongoing research and development to enhance the reliability and accuracy of AIGC models, ensuring they can be effectively used in real-world scenarios.
GALIP DFGAN SD_V14 SD_V15 Versatile SD_V21 SD_XL DALL·E 2
T2I 72.02 87.43 94.44 94.29 43.72 84.42 30.41 69.74
I2I - - 68.62 68.33 70.73 68.01 67.73 53.42
TI2I - - 62.84 62.35 62.28 58.56 29.89 -
Table 4: The results of Classification Mismatch Rate (MR).
GALIP DFGAN SD_V14 SD_V15 Versatile SD_V21 SD_XL DALL·E 2
T2I 0.23 0.23 0.25 0.25 0.26 0.25 0.26 0.26
I2I - - 0.38 0.38 0.69 0.50 0.82 0.42
TI2I - - 0.43 0.44 0.30 0.53 0.82 -
Table 5: The results of Segmentation IoU.
@@
In semantic segmentation, we evaluated the performance using the mean Intersection over Union (IoU) metric and show the results in Table 5. The mean IoU values for AIGC images varied between 0.23 and 0.82, indicating a wide range of segmentation accuracy. While some generated images approached the performance of natural images, a significant portion of AIGIs fell short, demonstrating poor segmentation results. This variability suggests that while AIGC models can sometimes produce high-quality segmentation, they often struggle to consistently match the accuracy of natural images.
@@
To further evaluate the performance of AI-generated images and natural images in downstream tasks, we conducted tests on the VQA model BLIP-2 (Li et al. 2023b). The evaluation involved three different types of questions: (1) identifying objects in the image (Object), (2) determining whether the image aligns with a given text by answering ”yes” or ”no” (Align), and (3) predicting the similarity probability between the image and the text (Prob.). Table 6 summarizes the discrepancies (%) between natural and AI-generated images across various models. For Object questions, the highest discrepancies occurred in the I2I setting, with SD_V14 and SD_V15 reaching 77.27% and 75.51%, respectively, while T2I discrepancies ranged from 36.08% (DALL·E 2) to 68.96% (DFGAN). Align questions showed lower discrepancies, with T2I values ranging from 24.72% (Versatile) to 34.65% (DFGAN) and I2I values from 15.90% (Versatile) to 29.24% (SD_V15). For similarity probability (Prob.), T2I discrepancies varied between 36.61% (DALL·E) and 62.01% (DFGAN), while I2I results ranged from 31.03% (DALL·E) to 62.80% (SD_V14), highlighting model-dependent variations.
These results indicate that AI-generated images, especially those produced under certain generative settings, face significant challenges in downstream VQA tasks compared to natural images. The findings also reveal performance variations across generative models, underscoring the need for improved alignment and semantic understanding in AI-generated images.
GALIP DFGAN SD_V14 SD_V15 Versatile SD_V21 SD_XL DALL·E 2
Object T2I 47.84 68.96 39.70 38.97 43.46 38.50 38.42 36.08
I2I - - 77.27 75.51 28.78 71.59 28.37 26.03
TI2I - - 41.18 38.06 37.64 39.27 26.69 -
Align T2I 29.81 34.65 27.52 28.03 24.72 31.99 29.78 29.79
I2I - - 28.48 29.24 15.90 27.07 16.33 25.50
TI2I - - 27.52 25.79 19.26 27.31 15.85 -
Prob. T2I 54.61 62.01 54.74 52.69 49.67 54.84 58.68 36.61
I2I - - 62.80 58.39 36.16 56.14 42.88 31.03
TI2I - - 52.75 48.71 43.69 46.30 43.24 -
Table 6: The results of VQA task w.r.t Object, Alignment, and Similarity probability.
SD_V21 SD_XL Versatile DALL·E 2

### v2

@@ 2. Semantic Alignment Discrepancy:
Generative models require the referenced prompts as semantic guidance to generate images. Therefore, semantic alignment can be an important quality indicator of the AIGIs. We use the widely used CLIP model (Radford et al., 2021a) to calculate the difference rates of Clip Scores, $DR_{CS}$ to represent alignment discrepancies between AIGIs and natural images, and provide a fine-grained analysis with respect to both unimodal and multimodal guidance.
3. Aesthetic Appeal Discrepancy:
@@
Aesthetic Appeal is to estimate the visual appeal and artistic quality of images, reflecting the visual attractiveness and artistic quality of images. We utilize the classical aesthetic measures NIMA (Esfandarani & Milanfar, 2018), LAION-AES (Schuhmann et al., 2022a), and QAlign (Wu et al., 2024b) as the quantitative metrics and compare the difference rates of aesthetic scores $DR_{aes}$ for the paired AI-generated and natural images.
4. Downstream Applicability Discrepancy:
@@
This aspect is devised to investigate the practical utility of AI-generated images in downstream tasks and evaluate whether AI-generated images can have different practical utilities with realistic natural images in downstream application tasks. We mainly focus on three common downstream tasks, e.g., image recognition, object segmentation and Multimodal Visual Question Answering (VQA). For the image recognition task, we evaluate the Difference recognition rates ($DR_{re}$) between AI-generated images and natural images when using a pre-trained image recognizer like ResNet-152 (He et al., 2016) model. For the object segmentation, we use the Intersection over Union (IoU) (Everingham et al., 2010) as quantitative measures and investigate the segmentation discrepancy for AI-generated images with U2Net (Qin et al., 2020). For VQA, we design three types of questions: ‘object’ (identifying objects in the image), ‘alignment’ (determining if the image aligns with a textual description, returning ‘yes’ or ‘no’), and ‘similarity probability’ (calculating the alignment probability between text and image). Using BLIP2 (Li et al., 2023b), we evaluate the Difference Rate ($DR$) between AI-generated images and reference natural images for each question type.
5. Human Validation:
@@
We involve human assessments to coordinate the above evaluation aspects. We develop the human assessment interface to demonstrate AI-generated images with the referenced natural images and text descriptions and collect human ratings (on a scale from 1 to 5) alongside the following three aspects: naive image quality, semantic alignment, and Aesthetic Appeal as human assessments. Human participants are provided with example images before the evaluation to understand high and low scores without being given specific numerical values, ensuring unbiased and informed ratings. Detailed information about human assessment procedures is provided in Appendix Sec. D.3.
@@
Through the integration of diverse and comprehensive evaluated dimensions, our D-Judge offers a systematic assessment solution to investigate and interpret the differences that still remain between AI-generated images and natural images.
@@
Figure 4: We evaluate nine image-generating models from 6 aspects. The numerical values in the radar chart represent the mean difference rate of each model.
4 Evaluation
@@
RQ1: What are the fine-grained naive image quality discrepancies between AIGIs and natural images?
@@
RQ2: How significant are the discrepancies in the semantic alignment of the given text with natural images and AIGIs in different types of guidance?
@@
4.1 Experimental Setting
We use the described quantitative evaluation metrics, which can be divided into Full-reference and No-reference, to present the $DR$ for involved assessment dimensions (more details of all metrics are presented in Appendix Sec. C.
@@ 4.1 Experimental Setting
For full-reference metrics, which are used for pixel-level image quality, visual content distribution, and downstream object segmentation task, i.e., IoU, we directly report the calculated value to reflect the difference cause it already can present the discrepancy.
@@
$\text{{DR}}=\frac{1}{n}\sum_{i=1}^{n}\frac{\lvert((M(\mathbf{X}_{i})-m_{\text{min}})-(M(\mathbf{N}_{i})-m_{\text{min}}))\rvert}{m_{\text{max}}-m_{\text{min}}}$ (1)
where $n$ is the number of images, $M$ represents the metric, $\mathbf{X}$ denotes the generated images, and $\mathbf{N}$ refers to the referenced natural images. Additionally, $m_{\text{min}}$ and $m_{\text{max}}$ represent the minimum and maximum values across all images, respectively, especially when the value range of the corresponding metric is undefined.
4.2 RQ1: Naive Quality Results
To evaluate the image quality of AI-generated images versus natural images, we conducted a comprehensive analysis using a variety of metrics that assess the discrepancies between AIGIs and natural images across pixel-level similarity, frame-level quality and visual content distribution. Our findings reveal substantial discrepancies between AIGIs and natural images across these dimensions.
4.2.1 RQ1.a Pixel-level
We evaluated the visual similarity between AI-generated and natural images using structural metrics like SSIM, LPIPS, DISTS, PSNR, etc. The mean similarity values of these metrics, as shown in Table 2, indicate that AI-generated images exhibit significantly lower similarity to natural counterparts. AI-generated images show a 30% to 65% reduction in SSIM, highlighting a major loss in structural fidelity. LPIPS and DISTS further reveal notable perceptual dissimilarities, with LPIPS scores ranging from 0.19 to 0.61. Additionally, PSNR values between 12.29 and 23.11 suggest that AI-generated images have higher noise levels, reducing their overall visual similarity.
Table 2: Full-reference Evaluation Results (I2I).
SD14 SD15 VD SD21 SDxl D·E 2
SSIM ↑ 0.35 0.35 0.46 0.43 0.70 0.37
PSNR ↑ 13.72 13.71 16.41 15.86 23.11 12.29
VIF ↑ 0.02 0.02 0.05 0.04 0.18 0.03
@@ 4.2 RQ1: Naive Quality Results
MAD ↓ 212.91 213.09 197.92 203.89 144.92 217.12
4.2.2 RQ1.b Frame-level
@@
We assessed the frame-level discrepancies of AI-generated and natural images using a suite of metrics, including DBCNN, HyperIQA, LIQE, PIQE, NRQM and QAlign. Figure 4 uses radar figures to visualize the mean difference rates from the six metrics, and indicates that there are significant quality differences between AI-generated and natural images at the frame level. These deviations are typically around 10%, irrespective of the guidance prompt used, with some instances showing quality differences as high as 40%. This underscores the considerable gaps between AIGIs and natural images in terms of global image quality. Moreover, the results of NRQM reveal that AI-generated images significantly differ from natural images in terms of naturalness. Specifically, AI-generated images show notable deviations from natural images in visual realism and adherence to natural scene statistics. The NRQM difference rates range from 5% to 32%, with images generated by D·E 3 exhibiting particularly high rates, underscoring the challenges AI-generated images encounter in replicating the inherent naturalness of real-world scenes. More detailed results with more metrics are presented in Appendix Sec. D.1 Figure 12.
@@
Figure 5: The FID (left, ↓) and Inception Score (right, ↑) of AI-generated images from different source models and guidance.
4.2.3 RQ1.c Visual Content Distribution
@@
We present the FID and Inception Score results, presented in Figure 5, suggesting that AI-generated images often exhibit relatively higher FID scores (up to 50) and lower Inception Scores (low to 14). This finding indicates that current generative models, particularly GAN-based models and earlier versions of Stable Diffusion, struggle to produce high-quality images as naturally collected images consistently. Notably, these models tend to generate better images when a reference image is provided (TI2I), as it has more detailed guidance.
@@
In conclusion, our evaluation of image quality across pixel-level, frame-level, and content distribution-level reveals significant disparities between AI-generated and natural images. These differences highlight the current limitations of AIGC technologies in matching the visual quality and realism of natural images.
4.3 RQ2: Alignment Results
@@
We calculate the difference rate of CLIP Scores ($DR_{CS}$) to assess how well the AIGIs align with natural ones in terms of semantic alignment across different guidance types: Text-to-Image (T2I), Image-to-Image (I2I), and Text-and-Image-to-Image (TI2I). And we find that AIGIs often struggle to maintain semantic alignment as natural ones, especially the image-only reference provided. The results, shown in Table 3, indicate that the $DR_{CS}$ range from 4.02% to 36.69%. We also find that even if the AIGIs can get a higher CS sometimes, especially with text and image-guided, they are looking stranger to humans. This suggests that current AIGC models often struggle to maintain high fidelity to the provided references, leading to variations that may impact the intended alignment. Our findings underscore the challenges that these models face in consistently generating content to accurately reflect the specified input conditions, highlighting both their strengths and limitations in real-world applications.
Table 3: The results of $DR_{CS}$ (%).
GALIP DFGAN SD14 SD15 VD SD21 SDxl D·E 2 D·E 3
T2I 12.16 16.11 10.67 10.53 9.71 10.90 11.59 10.93 10.19
I2I - - 33.43 33.57 4.89 24.04 3.56 5.53 -
TI2I - - 9.90 9.73 8.22 9.20 3.57 - -
4.4 RQ3: Aesthetic Appeal
@@
We assessed the aesthetic appeal of AI-generated and natural images using NIMA and LAION-AES metrics and found that AI-generated images generally fall short of natural images in terms of aesthetic quality. Figure 6 presents the aesthetic difference rate ($DR_{aes}$) results, which reveal that AI-generated images have noteworthy discrepancies compared to natural ones, with $DR_{aes}$ ranging from 3.51% to 27.66%. These findings indicate that while AI models can produce visually appealing images, AIGIs’ often lack the nuanced artistic quality and emotional impact that are characteristic of natural images.
Figure 6: The $DR_{aes}$ between AIGIs and natural ones.
4.5 RQ4: Applicability Results
@@
When evaluating the applicability of AI-generated images, we focused on three crucial downstream tasks: image recognition, semantic segmentation and VQA. Our findings indicate significant discrepancies between the performance of AIGC images and natural images in these tasks.
@@
We measured the classification Difference Rate (DR) (%) to identify the differences in predicted labels between AIGC images and their natural counterparts. The results in Table 8 showed a substantial gap, with the DR ranging from 27.46% to 94.29%, indicating that the generated images can’t accurately present the object as expected.
@@
We evaluated the performance using the Intersection over Union (IoU) metric and show the discrepancy results in Table 9. The $IoU$ values (higher is better) for AIGC images varied between 0.33 and 0.82, indicating a wide range of segmentation between AIGC images and their natural counterparts.
@@
Further, we apply these AI-generated image to the VQA task by BLIP-2 (Li et al., 2023b) involved three different types of questions: Object, Alignment (Align), and similarity probability (Prob.). Table 4 summarizes the discrepancies (%) between natural and AI-generated images across various questions. For Object, the highest discrepancies occurred in the I2I setting, with SD14 and SD15 reaching 96.41% and 96.95%, respectively, while T2I discrepancies ranged from 64.07% (DALL·E 3) to 88.43% (DFGAN). Align questions showed lower discrepancies, with T2I values ranging from 8.55% (DFGAN) to 21.18% (DALL·E 3) and I2I values from 7.87% (VD) to 10.07% (SD15). For similarity probability (Prob.), T2I discrepancies varied between 37.31% (DALL·E 2) and 62.69% (DFGAN), while I2I results ranged from 32.81% (DALL·E) to 61.56% (SD15).
Overall, our findings highlight significant limitations of AI-generated images in tasks such as image recognition, semantic segmentation, and visual question answering (VQA). While AI-generated images perform comparably to natural images in simpler tasks, such as binary text-image alignment (e.g., responding with “yes” or “no”), they fall short in more complex scenarios, including accurately estimating text-image alignment probabilities, identifying specific objects, and segmenting object areas. These shortcomings limit their broader applicability in diverse downstream tasks compared to natural images. Our results underscore the need for continued research and development to improve the reliability and accuracy of AIGC models, enabling their effective deployment in real-world applications.
Table 4: The Difference Rate (DR) of VQA tasks.
GALIP DFGAN SD14 SD15 VD SD21 SDxl D·E 2 D·E 3
Object T2I 73.34 88.43 68.72 68.63 74.21 68.34 66.93 57.85 64.07
I2I - - 96.41 96.95 50.43 91.14 51.65 46.29 -
TI2I - - 67.22 66.65 66.29 67.22 49.20 - -
Align T2I 10.57 8.55 14.02 15.79 12.32 13.02 13.00 12.99 21.18
I2I - - 9.71 9.53 7.87 10.07 8.01 8.93 -
TI2I - - 13.31 12.70 10.03 10.99 8.46 - -
Prob. T2I 53.13 62.69 53.07 54.21 50.57 54.95 57.09 37.31 49.22
I2I - - 61.05 61.56 39.57 58.67 43.64 32.81 -
TI2I - - 50.96 50.34 45.93 49.97 41.48 - -
4.6 Human Validation
To rigorously assess the differences between AI-Natural images and to evaluate the effectiveness of existing metrics in measuring AIGC image quality, we conducted human evaluations focusing on three key aspects: Image Quality, Alignment, and Aesthetic Appeal.
  • Image Quality: Participants assessed the overall quality of the images, considering factors such as clarity, detail, and the presence of artifacts.
  • Alignment: This aspect measured how well the generated images matched the given prompts.
  • Aesthetic: The aesthetic appeal of the images was rated based on visual attractiveness, composition, and overall artistic quality.
We excluded results for Applicability since this dimension inherently involves human-generated ground truth labels. Participants rated images on a scale from 1 (poor) to 5 (excellent). Our human evaluation results are shown in Table 5 and Figure 7 reveal significant discrepancies between AI-generated images and natural images across all three assessed aspects: quality, alignment, and aesthetic appeal. AIGC images consistently received lower scores across all three dimensions; AI-generated images generally received lower scores, often between 2 and 3, compared to natural images. This emphasizes the need for further advancements to narrow these gaps. These findings underscore the importance of continued research and development in AI image generation to bring AIGC images closer to the standards of natural images.
Moreover, the result in Figure 7 from human validation indicates that traditional quantitative metrics often fail to align with human preferences, suggesting that they are not fully compatible with assessing AIGC images. This highlights the necessity for developing new measures tailored specifically to the unique characteristics of AI-generated content.
Table 5: Human score results across three aspects.
SD21 SDxl VD D·E 2